OpenStudy (anonymous):

consider F(x) = 1-e^x. Find the slope of the graph of f(x) at the point where the graph crosses the x-axis.

OpenStudy (chaise):

I might be wrong but is the slope -1? Can anyone else concur?

OpenStudy (anonymous):

huh, how did you get that? :O

OpenStudy (chaise):

Differentiate the equation and substitute, you get -e^x. When the graph cuts the x axis, x is equal to zero. -e^x =-e^0 =-1

OpenStudy (anonymous):

To clarify, you use both F(x) and f(x) in your question, by f(x) do you mean the derivative of F(x), or did you mean them both to be the same?

OpenStudy (chaise):

I'm not sure if this is right, your question is a bit funny, which is what nbouscal is clarifying now.

OpenStudy (anonymous):

oh, I see, thank's chaise! there was a follow-up. could you help with that as well? the question's this: Find the equation of the tangent line to the curve at this point.

OpenStudy (chaise):

well, you know the slope of the equation is -1, and if the tangent is at the point where the original graph cuts the x axis, then the y intercept of the equation of the tangent will also be 0. The equation of the line is: y=-1x+0 The y intercept is 0.